# Robust Global Fréchet Regression via Weight Regularization

### v1

@@ Abstract
In recent years, the regression methods for response variables on manifolds have become increasingly popular, including probability distribution responses (hartung2001tests), covariance matrices (newey1986simple), network responses (bar2004response), and other complex objects. Their use is becoming more widespread in real-world data analysis, particularly in medicine, geological science, and logistics, as (marron2014overview). However, traditional regression techniques, which are designed for Euclidean-valued responses, are inadequate for modeling such complex data structures. To address this challenge, several recent studies have investigated regression models for non-Euclidean and manifold-valued data. fletcher2011geodesic proposed a geodesic regression model called “global Fréchet regression”, which is the natural generalization of linear regression and is parameterized by an intercept and slope term. miller2004computational and jupp1987fitting proposed an unrolling method on shape spaces. Fréchet regression, based on the Fréchet mean, has emerged as a powerful extension, enabling regression analysis when the response variable lies in a non-Euclidean metric space.
@@
A notable limitation of the existing Fréchet regression is the sensitivity to outlying observations. However, research on robust regression methods in manifold spaces remains very limited. To the best of our knowledge, the only related work is that of lee2024huber and NIPS2009_92977ae4, who proposed and systematically analysed the Huber mean on Riemannian manifolds and provided an iterative algorithm for parameter estimation. It should be noted that each iteration of this algorithm requires geometric operations on the manifold, such as the exponential and logarithmic maps. Specifically, the logarithmic map takes each data point. It transforms it to a vector in the tangent space at the current mean, effectively describing the direction and distance from the current mean to that data point on the manifold. Since these log and exp maps do not have explicit analytical formulas for most manifolds and must be computed numerically, the computational cost of each iteration is substantially increased.
@@
In this work, we propose a novel approach to the global robust Fréchet regression developed by petersen2019frechet. The original Fréchet regression provides a principled approach for modelling regression relationships between vectors of real-valued predictors and complex response objects residing in a general metric space. However, similar to the aforementioned method, the standard Fréchet regression lacks robustness, making it sensitive to outliers and deviations. Thus, we incorporate a weight parameter (taking values on $[0,1]$) into the original objective function of the Fréchet regression and give the Elastic net penalty term to the weight paraemter. Under this framework, observations identified as outliers are assigned weights close to zero, effectively reducing their influence on the regression estimation, whereas typical observations receive weights near one. However, direct regularization of the weight parameters themselves would undesirably shrink all weights towards zero, thereby diminishing the influence of all observations, including those that are not outliers. Our new methodology offers two key advantages. First, under both the Frobenius distance and the $L_{2}$ Wasserstein distance, it allows for closed-form solutions for the estimators, thereby facilitating efficient computation. Second, our simulation demonstrates that the proposed algorithm exhibits rapid convergence, often requiring only a small number of iterations to achieve stable estimates. For the selection of optimal tuning parameters in the regularization term, we adopt the Bayesian information criterion (BIC), following gao2016penalized, who proposed a weighted model for response variables in the Euclidean space. Building upon this approach, we extend the methodology to accommodate situations where the response variables reside in non-Euclidean spaces.
@@
The rest of the paper is organized as follows. Section 2 provides an overview of global robust Fréchet regression, introduces the framework of robust Fréchet regression with weight regularization, discusses the linear convergence properties of the proposed optimization algorithm, and describes the procedure for selecting tuning parameters using the BIC criterion. In Section 3, we demonstrate the applicability of our approach to both matrix-valued and distribution-valued responses,and present a fixed-point algorithm for implementation, and report simulation results along with analysis on real-world datasets to illustrate the effectiveness of the proposed method. In Section 4, we provide a brief discussion of the methods and possible extensions. R code implementing the proposed method is available at the GitHub repository (https://github.com/lee1995hao/robust-FR).
@@
$\bar{Y}={\rm argmin}_{u}\sum_{i=1}^{n}d(Y_{i},u)^{2},$
where $d(\cdot,\cdot)$ is a distance. The existence of $\bar{Y}$ is always guaranteed, although uniqueness depends on the curvature properties of the metric space (e.g., Hilbert spaces or non-positively curved spaces ensure uniqueness). When the associated (Euclidean) covariate $X_{i}$ is available, the global Fréchet regression function (petersen2019frechet) can be obtained by
$m(x)={\rm argmin}_{u}\sum_{i=1}^{n}g(X_{i},x)d(Y_{i},u)^{2},$ (1)
@@
where $W_{i}>0$ is a weight parameter. Here $w=\{W_{1},\ldots,W_{n}\}$ represents a set of weights, and the weight $W_{i}$ plays a critical role in determining the contribution of each observation to the loss function. Specifically, when $W_{i}=1$, the corresponding observation $Y_{i}$ is fully utilized in the estimation process, whereas if $W_{i}=0$, the information from $Y_{i}$ is entirely excluded. Initially, the weights $W_{i}$ should adaptively reflect the outlyingness of each observation, such that $W_{i}=1$ for genuine (non-outlying) observations and $W_{i}=0$ for outliers. Since the classification of observations as outlying or non-outlying is unknown a priori, $W_{i}$ is treated as an unknown parameter to be jointly estimated alongside the regression function $u$.
While $w$ is a high-dimensional parameter, we can assume sparsity for $w$ in the sense that most elements in $w$ are $1$, indicating that most observations are genuine (non-outlying) observations. Hence, in the estimation of $w$, we introduce a regularization term, where a similar approach is typically adopted in the estimation of Shift in the robust regression (she2011outlier). Specifically, we employ the Elastic net penalty (zou2005regularization) for $1-W_{i}$. This penalty simultaneously enforces sparsity and smoothness in the estimation of $W_{i}$, facilitating effective differentiation between outliers and non-outlying observations. We therefore define the robust Fréchet regression function with the following objective function:
@@
where $\lambda$ and $\gamma$ are tuning parameters. Then, the regression function and weight parameter can be obtained as $(\hat{u},\hat{w})={\rm argmin}_{u,w\in[0,1]^{n}}Q(u,w)$. This optimization problem can be easily solved by an iterative algorithm described in Section 2.4. Given the regression function $u$, the optimal weight minimizing (3) can be obtained as follows:
@@
The optimal weight $({\widetilde{W}}_{1}(u),\ldots,{\widetilde{W}}_{n}(u))={\rm argmin}_{u\in[0,1]^{n}}Q(u,w)$ is obtained as
${\widetilde{W}}_{i}(u)=\begin{cases}1,&g(X_{i},x)d^{2}(Y_{i},u)\in[0,\lambda]\\ 1-\dfrac{1}{2\gamma}\bigl\{g(X_{i},x)d^{2}(Y_{i},u)-\lambda\bigr\},&g(X_{i},x)d^{2}(Y_{i},u)\in(\lambda,\lambda+2\gamma)\\ 0,&g(X_{i},x)d^{2}(Y_{i},u)\in[\lambda+2\gamma,\infty)\end{cases}$ (4)
@@
In Figure 3, we present the shape of the adaptive weight ${\widetilde{W}}_{i}(u)$ as a function of the weighted distance $g(X_{i},x)d^{2}(Y_{i},u)$ under four cases of $(\lambda,\gamma)$. The curve of $W_{i}$ decreases as $d^{2}(Y_{i},u)$ increases and $\lambda>0$ and $\gamma>0$ as Figure 3 . The Tuning parameter $\lambda$ establishes the threshold for $W_{i}=1$, controlling the quantity of total normal values. Conversely, for a fixed value of $\lambda$, the Tuning parameter $\gamma$ determines the threshold for $W_{i}=0$, which controls the quantity of outlier values.
@@
where $B(u)=\sum_{i=1}^{n}\lambda|1-{\widetilde{W}}_{i}(u)|+\gamma\{1-{\widetilde{W}}_{i}(u)\}^{2}$. The first term of $\widetilde{Q}(u)$ can be regarded as a version of weighted least squares (kiers1997weighted) for the global Fréchet regression with robust weight ${\widetilde{W}}_{i}(u)$. Regarding the second term $B(u)$, we let $\omega$ be a proportion of outlying observations such that $n^{-1}\sum_{i=1}^{n}I\{{\widetilde{W}}_{i}(u_{\ast})<1\}=\omega\in(0,1)$ for the true regression function $u_{\ast}$. Then, it holds that $0\leq B(u_{\ast})\leq\omega n$, indicating that the second term could be negligible when $\omega$ is relatively small (e.g. $\omega=0.05$). Hence, the profiled objective function $\widetilde{Q}(u)$ is approximately equal to the outcome-dependent weighted loss function around the neighborhood of the true regression function $u_{\ast}$. A notable feature of the joint minimization of $Q(u,w)$ is that the optimization can be efficiently performed through an iterative algorithm.
@@
We here discuss the necessity of the Elastic net penalty in (3). Following the regularized estimation of Shift value, one may consider the following $L_{1}$-penalized objective function:
@@
for $W_{i}>0$, instead of the Elastic net penalty given in (3). Note that the objective function $Q^{\dagger}(u,w)$ is equivalent to (3) with $\gamma=0$ (without quadratic term).
@@
${\widetilde{W}}_{i}^{\dagger}(u)=\begin{cases}1,&g(X_{i},x)d^{2}(Y_{i},u)\in[0,\lambda)\\ [0,1],&g(X_{i},x)d^{2}(Y_{i},u)=\lambda\\ 0,&g(X_{i},x)d^{2}(Y_{i},u)\in(\lambda,\infty).\end{cases}$
@@
$u^{(s+1)}(x)\ \leftarrow\ {\rm argmin}_{u}\sum_{i=1}^{n}W_{i}^{(s+1)}g(X_{i},x)d^{2}(y_{i},u).$
@@
    There exists a constants, $D_{u}>0$, such that $d(Y_{i},u)\leq D_{u}$ for all $i$ and $u\in\mathcal{U}$.
@@
    $g(X_{i},x)<\infty$ for all $x\in\mathcal{X}$.
@@
    For $w=(W_{1},\ldots,W_{n})\in[0,1]^{n}$, define a map $\Phi(w)$ as
    $\Phi(w)={\rm argmin}_{u\in\mathcal{U}}\sum_{i=1}^{n}W_{i}g(X_{i},x)d^{2}(Y_{i},u).$
    There exists a constant $C_{u}>0$ such that $d(\Phi(w_{1}),\Phi(w_{2}))\leq C_{u}\|w_{1}-w_{2}\|$ for all $w_{1},w_{2}\in[0,1]^{n}$.
The conditions (C1) and (C2) are finiteness of the metric space and weight values. The conditions (C3) and (C4) are the Lipschitz conditions for the function $d^{2}(Y_{i},\cdot)$ and the updating function $\Phi(w)$ for $u$ given $w$. Then, we have the linear convergence property of Algorithm 1.
@@
Under regularity conditions (C1)-(C4) , $d(u^{(s)},u^{\ast})\leq\rho^{s}d(u^{(0)},u^{\ast})$, for the minimizer $u^{\ast}$, so that Algorithm 1 exhibits linear convergence if $\rho<1$.
@@
There are two tuning parameters, $(\lambda,\gamma)$, in the objective function (3), which would significantly control the downweighting of outliers as in (4). Here, we propose a data-dependent method for selecting the tuning parameters. Let $\hat{W}_{i}(\lambda,\gamma)$ and $\hat{u}(X_{i};\lambda,\gamma)$ be estimates minimizing (3) with fixed $(\lambda,\gamma)$. Then, the square of “residual” can be defined as $d^{2}\{Y_{i},\hat{u}(X_{i};\lambda,\gamma)\}$. Based on this quantity, we employ the following Bayesian information criterion (gao2016penalized):
@@
where $\hat{k}(\lambda,\gamma)=\sum_{i=1}^{n}{\rm I}\{\hat{W}_{i}(\lambda,\gamma)<1\}$ is the number of “outliers” under the tuning parameter $(\lambda,\gamma)$. A similar criterion was introduced in she2011outlier. The BIC formula (5) indicates a trade-off between the goodness of fit and the number of suspected outliers. In fact, the first term in (5) measures the goodness of fit while the second term measures the model robustness.
The optimal tuning parameter can be defined as the minimizer of the criterion (5). However, according to she2011outlier, when the selected values of $\lambda$ and $\gamma$ result in a huge number of estimated outliers, it is often observed that the discriminative power of BIC substantially deteriorates. Therefore, it is recommended to define the lower bounds of $\lambda$ and $\gamma$ as the values corresponding to when the number of outliers exceeds 30% of the total sample size, and the upper bounds as the maximum values ensuring that all data points are classified as non-outliers. Within this bounded interval, parameter selection and model screening based on BIC should be conducted to enhance the robustness and accuracy of outlier detection.
@@
When $Y_{i}$ is a matrix or network, one may use the Frobenius metric (hitchin1997frobenius) defined as $d(L_{1},L_{2})=\sqrt{\text{tr}\left[(L_{1}-L_{2})^{\top}(L_{1}-L_{2})\right]}$ for some matrices $L_{1}$ and $L_{2}$. In this case, the updating step for $u$ given $w$ is equivalent to minimizing
@@
When $Y_{i}$ is a distribution, $L_{2}$-Wasserstein distance can be employed to quantify the differences between two distributions. Regarding The $L_{2}$-Wasserstein distance (e.g. panaretos2016amplitude; turner2014frechet) between two distributions $F_{1},F_{2}$ can be defined as $d(F_{1},F_{2})=\|F_{1}^{-1}(z)-F_{2}^{-1}(z)\|_{2}$, where $F_{1}^{-1}(z)$ and $F_{2}^{-1}(z)$ represent the quantile functions for $z\in[0,1]$ and $\|\cdot\|_{2}$ denotes $L^{2}$-norm. Under the settings, the updating step for $u$ given $w$ is
$\sum_{i=1}^{n}W_{i}g(X_{i},x)\|F_{i}^{-1}(z)-u\|_{2}^{2},$
@@
    We generate $q$-dimensional covariate $X$ from the uniform distribution on $0,1]^{p}$, and the response matrix $Y$ is generated via symmetric matrix variate normal distribution, following qiu2024random. The $(j,k)$-element of $Y$, denoted by $Y_{jk}$, is defined as $Y_{jk}=\exp\{0.2Z_{jk}+D_{jk}(X)\}$, where $Z_{jk}\sim N(0,1)$ for $j=k$ and $Z_{jk}\sim N(0,1/2)$ for $j\neq k$. Here $D_{jk}(X)=1$ for $j=k$ and $D_{jk}(X)=U_{jk}\cos(4\pi(\beta^{T}X))$ with $U_{jk}\sim U(0,1)$, where $\beta=(0.1,\,0.2,\,0.3,\,0.4,\,0.5,\,0,\ldots,0)^{\top}$.
@@
Table 1 summarizes the average MSE values averaged over 100 Monte Carlo replications of the standard and robust Fréchet regression estimators under two scenarios of DGP and five contamination settings. The Monte Carlo standard errors are reported in parentheses. Under DGP (I) without contamination, both estimators exhibit nearly identical MSEs, indicating that the robust modification preserves efficiency in the uncontaminated setting. As the contamination proportion increases, the MSE of the non-robust method rises sharply, whereas the robust Fréchet regression method displays only a mild increase. This tendency holds for both $n=50$ and $n=100$. The corresponding relative MSEs (standard over robust) exceed 20 in the highest contamination scenarios. In DGP (II), the original Fréchet regression method has a modest advantage in the uncontaminated setting, while the original Fréchet regression exhibits a much deeper escalation of MSE as the contamination ratio increases than the robust Fréchet regression. We further report the BIC-selected tuning parameters $\lambda$ and $\gamma$ value across data scenarios of DGP (I); the results are presented in the Table 2. We find that, as the contamination proportion and Shift value increase, the BIC-selected $\lambda$ and $\gamma$ also increase. These results suggest that $\gamma$ also plays a critical role in downweighting the influence of heavily contaminated and high-bias observations.
@@
Table 2: The value of $\lambda$ and $\gamma$ for matrix response under various data configurations (sample size $n$ and other parameter settings). Values are mu;multiplied by $10^{2}$.
@@
Figure 2: True networks(first), robust fitted networks(second), network regression method proposed by zhou2022network(third), and non-robust fitted networks (fourth) on May 16, 2020, corresponds to the day when the number of new COVID-19 cases was 134.
Figure 3: The absolute error heat map between the true network value, the network regression method proposed by zhou2022network, and the predicted network value on May 16, 2020, corresponds to the day when the number of new COVID-19 cases was 134.

### v2

@@ Introduction
In recent years, the regression methods for response variables on manifolds have become increasingly popular, including probability distribution responses (Hartung and Knapp, 2001), covariance matrices (Newey and West, 1986), network responses (Bar-Yam and Epstein, 2004), and other complex objects. Their use is becoming more widespread in real-world data analysis, particularly in medicine, geological science, and logistics, as (Marron and Alonso, 2014). However, traditional regression techniques, which are designed for Euclidean-valued responses, are inadequate for modeling such complex data structures. To address this challenge, several recent studies have investigated regression models for non-Euclidean and manifold-valued data. Fletcher (2011) proposed a geodesic regression model called “global Fréchet regression”, which is the natural generalization of linear regression and is parameterized by an intercept and slope term. Miller (2004) and Jupp and Kent (1987) proposed an unrolling method on shape spaces. Fréchet regression, based on the Fréchet mean, has emerged as a powerful extension, enabling regression analysis when the response variable lies in a non-Euclidean metric space.
@@
A notable limitation of the existing Fréchet regression is the sensitivity to outlying observations. However, research on robust regression methods in manifold spaces remains very limited. To the best of our knowledge, the only related work is that of Lee and Jung (2024) and Hein (2009), who proposed and systematically analysed the Huber mean on Riemannian manifolds and provided an iterative algorithm for parameter estimation. It should be noted that each iteration of this algorithm requires geometric operations on the manifold, such as the exponential and logarithmic maps. Specifically, the logarithmic map takes each data point. It transforms it to a vector in the tangent space at the current mean, effectively describing the direction and distance from the current mean to that data point on the manifold. Since these log and exp maps do not have explicit analytical formulas for most manifolds and must be computed numerically, the computational cost of each iteration is substantially increased.
@@
In this work, we propose a novel approach to the global robust Fréchet regression developed by Petersen and Müller (2019). The original Fréchet regression provides a principled approach for modeling regression relationships between vectors of real-valued predictors and complex response objects residing in a general metric space. However, similar to the aforementioned method, the standard Fréchet regression lacks robustness, making it sensitive to outliers and deviations. Thus, we incorporate a weight parameter (taking values on $[0,1]$) into the original objective function of the Fréchet regression and give the Elastic net penalty term to the weight parameter. Under this framework, observations identified as outliers are assigned weights close to zero, effectively reducing their influence on the regression estimation, whereas typical observations receive weights near one. However, direct regularization of the weight parameters themselves would undesirably shrink all weights towards zero, thereby diminishing the influence of all observations, including those that are not outliers. Our new methodology offers two key advantages. First, under both the Frobenius distance and the $L_{2}$ Wasserstein distance, it allows for closed-form solutions for the estimators, thereby facilitating efficient computation. Second, our simulation demonstrates that the proposed algorithm exhibits rapid convergence, often requiring only a small number of iterations to achieve stable estimates. For the selection of optimal tuning parameters in the regularization term, we adopt the Bayesian information criterion (BIC), following Gao and Fang (2016), who proposed a weighted model for response variables in the Euclidean space. Building upon this approach, we extend the methodology to accommodate situations where the response variables reside in non-Euclidean spaces.
@@
The rest of the paper is organized as follows. Section 2 provides an overview of global robust Fréchet regression, introduces the framework of robust Fréchet regression with weight regularization, discusses the linear convergence properties of the proposed optimization algorithm, and describes the procedure for selecting tuning parameters using the BIC criterion. In Section 3, we demonstrate the applicability of our approach to both matrix-valued and distribution-valued responses,and present a fixed-point algorithm for implementation, and report simulation results along with analysis on real-world datasets to illustrate the effectiveness of the proposed method. In Section 4, we provide a brief discussion of the methods and possible extensions. R code implementing the proposed method is available at the GitHub repository (https://github.com/lee199950120HAO/robust-FR).
@@
$\bar{Y}=\mathop{\rm argmin}_{u\in\mathcal{U}}\sum_{i=1}^{n}d(Y_{i},u)^{2},$
where $d(\cdot,\cdot)$ is a distance. The existence of $\bar{Y}$ is always guaranteed, although uniqueness depends on the curvature properties of the metric space (e.g., Hilbert spaces or non-positively curved spaces ensure uniqueness). When the associated (Euclidean) covariate $X_{i}$ is available, the global Fréchet regression function (Petersen and Müller, 2019) can be obtained by
$m(x)=\mathop{\rm argmin}_{u\in\mathcal{U}}\sum_{i=1}^{n}g(X_{i},x)d(Y_{i},u)^{2},$ (1)
@@
where $1\geq W_{i}\geq 0$ is a weight parameter. Here $w=\{W_{1},\ldots,W_{n}\}$ represents a set of weights, and the weight $W_{i}$ plays a critical role in determining the contribution of each observation to the loss function. Specifically, when $W_{i}=1$, the corresponding observation $Y_{i}$ is fully utilized in the estimation process, whereas if $W_{i}=0$, the information from $Y_{i}$ is entirely excluded. Initially, the weights $W_{i}$ should adaptively reflect the outlyingness of each observation, such that $W_{i}=1$ for genuine (non-outlying) observations and $W_{i}=0$ for outliers. Since the classification of observations as outlying or non-outlying is unknown a priori, $W_{i}$ is treated as an unknown parameter to be jointly estimated alongside the regression function $u$.
While $w$ is a high-dimensional parameter, we can assume sparsity for $w$ in the sense that most elements in $w$ are $1$, indicating that most observations are genuine (non-outlying) observations. Hence, in the estimation of $w$, we introduce a regularization term, where a similar approach is typically adopted in the estimation of Shift in the robust regression (She and Owen, 2011). Specifically, we employ the Elastic net penalty (Zou and Hastie, 2005) for $1-W_{i}$. This penalty simultaneously enforces sparsity and smoothness in the estimation of $W_{i}$, facilitating effective differentiation between outliers and non-outlying observations. We therefore define the robust Fréchet regression function with the following objective function:
@@
where $\lambda$ and $\gamma$ are tuning parameters. Then, the regression function and weight parameter can be obtained as $(\hat{u},\hat{w})={\rm argmin}_{u\in\mathcal{U},w\in[0,1]^{n}}Q(u,w)$. This optimization problem can be easily solved by an iterative algorithm described in Section 2.4. Given the regression function $u$, the optimal weight minimizing (3) can be obtained as follows:
@@
The optimal weight $({\widetilde{W}}_{1}(u),\ldots,{\widetilde{W}}_{n}(u))={\rm argmin}_{w\in[0,1]^{n}}Q(u,w)$ is obtained as
${\widetilde{W}}_{i}(u)=\begin{cases}1,&g(X_{i},x)d^{2}(Y_{i},u)\in[-\infty,\lambda]\\ 1-\dfrac{1}{2\gamma}\bigl\{g(X_{i},x)d^{2}(Y_{i},u)-\lambda\bigr\},&g(X_{i},x)d^{2}(Y_{i},u)\in(\lambda,\lambda+2\gamma)\\ 0,&g(X_{i},x)d^{2}(Y_{i},u)\in[\lambda+2\gamma,\infty)\end{cases}$ (4)
@@
In Figure 3, we present the shape of the adaptive weight ${\widetilde{W}}_{i}(u)$ as a function of the weighted distance $g(X_{i},x)d^{2}(Y_{i},u)$ under four cases of $(\lambda,\gamma)$. The curve of $W_{i}$ decreases as $d^{2}(Y_{i},u)$ increases and $\lambda>0$ and $\gamma>0$ as Figure 3 . The tuning parameter $\lambda$ establishes the threshold for $W_{i}=1$, controlling the quantity of total normal values. Conversely, for a fixed value of $\lambda$, the Tuning parameter $\gamma$ determines the threshold for $W_{i}=0$, which controls the quantity of outlier values. Although $W_{i}$ equals 1 whenever $g(X_{i},x)<0$, the sub-Gaussian assumption imposed in the convergence analysis guarantees that $g(X_{i},x)$ remains bounded with exponential tails. As the result, the product $g(X_{i},x),d^{2}(Y_{i},u)$ cannot reach excessively a large value.
@@
where $B(u)=\sum_{i=1}^{n}\lambda|1-{\widetilde{W}}_{i}(u)|+\gamma\{1-{\widetilde{W}}_{i}(u)\}^{2}$. The first term of $\widetilde{Q}(u)$ can be regarded as a version of weighted least squares (Kiers, 1997) for the global Fréchet regression with robust weight ${\widetilde{W}}_{i}(u)$. Regarding the second term $B(u)$, we let $\omega$ be a proportion of outlying observations such that $n^{-1}\sum_{i=1}^{n}I\{{\widetilde{W}}_{i}(u_{\ast})<1\}=\omega\in(0,1)$ for the true regression function $u_{\ast}$. Then, it holds that $0\leq B(u_{\ast})\leq\omega n$, indicating that the second term could be negligible when $\omega$ is relatively small (e.g. $\omega=0.05$). Hence, the profiled objective function $\widetilde{Q}(u)$ is approximately equal to the outcome-dependent weighted loss function around the neighborhood of the true regression function $u_{\ast}$. A notable feature of the joint minimization of $Q(u,w)$ is that the optimization can be efficiently performed through an iterative algorithm.
@@
We here discuss the necessity of the Elastic net penalty in (3). Following the regularized estimation of the shift value, one may consider the following $L_{1}$-penalized objective function:
@@
instead of the Elastic net penalty given in (3). Note that the objective function $Q^{\dagger}(u,w)$ is equivalent to (3) with $\gamma=0$ (without quadratic term).
@@
${\widetilde{W}}_{i}^{\dagger}(u)=\begin{cases}1,&g(X_{i},x)d^{2}(Y_{i},u)\in(-\infty,\lambda)\\ [0,1],&g(X_{i},x)d^{2}(Y_{i},u)=\lambda\\ 0,&g(X_{i},x)d^{2}(Y_{i},u)\in(\lambda,\infty).\end{cases}$
@@
$u^{(s+1)}(x)\ \leftarrow\ \mathop{\rm argmin}_{u\in\mathcal{U}}\sum_{i=1}^{n}W_{i}^{(s+1)}g(X_{i},x)d^{2}(y_{i},u).$
@@
    There exists a constant, $D_{u}>0$, such that $d(Y_{i},u)\leq D_{u}$ for all $i$ and $u\in\mathcal{U}$.
@@
    There exists a constant, $D_{g}>0$, such that $|g(X_{i},x)|\leq D_{g}$ for all $i$ and $x\in\mathcal{X}$.
  • (C3)
    $w=(W_{1},\ldots,W_{n})\in\mathcal{W}=\{w\in[0,1]^{n}\mid\sum_{i=1}^{n}W_{i}\geq\xi n\}$ for some $\xi\in(0,1)$.
@@
    For $w\in\mathcal{W}$, define a map $\Phi(w)$ as
    $\Phi(w)=\mathop{\rm argmin}_{u\in\mathcal{U}}\sum_{i=1}^{n}W_{i}g(X_{i},x)d^{2}(Y_{i},u).$
    There exists a constant $C_{u}>0$ such that $d(\Phi(w_{1}),\Phi(w_{2}))\leq C_{u}\|w_{1}-w_{2}\|$ for all $w_{1},w_{2}\in\mathcal{W}$.
The conditions (C1) and (C2) are finiteness of the metric space and weight values. The condition (C3) is related to the maximum number of outliers. If the number of samples meeting $\sup_{x\in\mathcal{X},u\in\mathcal{U}}g(X_{i},x)d^{2}(Y_{i},u)<\lambda$, regarded as genuine observations, is larger than $\xi n$ then (C3) holds from (4). The conditions (C4) and (C5) are the Lipschitz conditions for the function $d^{2}(Y_{i},\cdot)$ and the updating function $\Phi(w)$ for $u$ given $w$. Then, we have the linear convergence property of Algorithm 1.
@@
Under regularity conditions (C1)-(C5), define $\rho=(2\gamma)^{-1}\sqrt{n}C_{u}L_{d}D_{g}$. Then $d(u^{(s)},u^{\ast})\leq\rho^{s}d(u^{(0)},u^{\ast})$, for the minimizer $u^{\ast}$, so that Algorithm 1 exhibits linear convergence if $\rho<1$.
@@
There are two tuning parameters $\lambda$ and $\gamma$ in the objective function (3), which would significantly control the downweighting of outliers as in (4). Here, we propose a data-dependent method for selecting the tuning parameters. Let $\hat{W}_{i}(\lambda,\gamma)$ and $\hat{u}(X_{i};\lambda,\gamma)$ be estimates minimizing (3) with fixed $(\lambda,\gamma)$. Then, the square of “residual” can be defined as $d^{2}\{Y_{i},\hat{u}(X_{i};\lambda,\gamma)\}$. Based on this quantity, we employ the following Bayesian information criterion (Gao and Fang, 2016):
@@
where $\hat{k}(\lambda,\gamma)=\sum_{i=1}^{n}{\rm I}\{\hat{W}_{i}(\lambda,\gamma)<1\}$ is the number of “outliers” under the tuning parameter $(\lambda,\gamma)$. A similar criterion was introduced in She and Owen (2011). The BIC formula (5) indicates a trade-off between the goodness of fit and the number of suspected outliers. In fact, the first term in (5) measures the goodness of fit while the second term measures the model robustness.
The optimal tuning parameter can be defined as the minimizer of the criterion (5). However, according to She and Owen (2011), when the selected values of $\lambda$ and $\gamma$ result in a huge number of estimated outliers, it is often observed that the discriminative power of BIC substantially deteriorates. Therefore, it is recommended to define the lower bounds of $\lambda$ and $\gamma$ as the values corresponding to when the number of outliers exceeds 30% of the total sample size, and the upper bounds as the maximum values ensuring that all data points are classified as non-outliers. Within this bounded interval, parameter selection and model screening based on BIC should be conducted to enhance the robustness and accuracy of outlier detection.
@@
When $Y_{i}$ is a matrix or network, one may use the Frobenius metric (Hitchin, 1997) defined as $d(L_{1},L_{2})=\sqrt{\text{tr}\left[(L_{1}-L_{2})^{\top}(L_{1}-L_{2})\right]}$ for some matrices $L_{1}$ and $L_{2}$. In this case, the updating step for $u$ given $w$ is equivalent to minimizing
@@
When $Y_{i}$ is a distribution, $L_{2}$-Wasserstein distance can be employed to quantify the differences between two distributions. Regarding the $L_{2}$-Wasserstein distance (e.g. Panaretos and Zemel, 2016; Turner et al., 2014) between two distribution functions $F_{1},F_{2}$ can be defined as $d(F_{1},F_{2})=\|F_{1}^{-1}-F_{2}^{-1}\|_{2}$, where $F_{1}^{-1}(z)$ and $F_{2}^{-1}(z)$ represent the quantile functions for $z\in[0,1]$ and $\|\cdot\|_{2}$ denotes $L^{2}$-norm. Under the settings, the updating step for $u$ given $w$ is
$\sum_{i=1}^{n}W_{i}g(X_{i},x)\|F_{i}^{-1}-u\|_{2}^{2},$
@@
    We generate $q$-dimensional covariate $X$ from the uniform distribution on $0,1]^{p}$, and the response matrix $Y$ is generated via symmetric matrix variate normal distribution, following Qiu et al. (2024). The $(j,k)$-element of $Y$, denoted by $Y_{jk}$, is defined as $Y_{jk}=\exp\{0.2Z_{jk}+D_{jk}(X)\}$, where $Z_{jk}\sim N(0,1)$ for $j=k$ and $Z_{jk}\sim N(0,1/2)$ for $j\neq k$. Here $D_{jk}(X)=1$ for $j=k$ and $D_{jk}(X)=U_{jk}\cos(4\pi(\beta^{T}X))$ with $U_{jk}\sim U(0,1)$, where $\beta=(0.1,\,0.2,\,0.3,\,0.4,\,0.5,\,0,\ldots,0)^{\top}$.
@@
Table 1 summarizes the average MSE values averaged over 100 Monte Carlo replications of the standard and robust Fréchet regression estimators under two scenarios of DGP and five contamination settings. The Monte Carlo standard errors are reported in parentheses. Under DGP (I) without contamination, both estimators exhibit nearly identical MSEs, indicating that the robust modification preserves efficiency in the uncontaminated setting. As the contamination proportion increases, the MSE of the non-robust method rises sharply, whereas the robust Fréchet regression method displays only a mild increase. This tendency holds for both $n=50$ and $n=100$. The corresponding relative MSEs (standard over robust) exceed 20 in the highest contamination scenarios. In DGP (II), the original Fréchet regression method has a modest advantage in the uncontaminated setting, while the original Fréchet regression exhibits a much deeper escalation of MSE as the contamination ratio increases than the robust Fréchet regression. We further report the BIC-selected tuning parameters $\lambda$ and $\gamma$ value across data scenarios of DGP (I); the results are presented in the Table 2. We find that, as the contamination proportion and the shift value increase, the BIC-selected $\lambda$ and $\gamma$ also increase. These results suggest that $\gamma$ also plays a critical role in downweighting the influence of heavily contaminated and high-bias observations.
@@
Table 2: The value of $\lambda$ and $\gamma$ for matrix response under various data configurations (sample size $n$ and other parameter settings). Values are multiplied by $10^{2}$.
@@
Figure 2: True networks(first), robust fitted networks(second), network regression method proposed by Zhou and Müller (2022)(third), and non-robust fitted networks (fourth) on May 16, 2020, corresponds to the day when the number of new COVID-19 cases was 134.
Figure 3: The absolute error heat map between the true network value, the network regression method proposed by Zhou and Müller (2022), and the predicted network value on May 16, 2020, corresponds to the day when the number of new COVID-19 cases was 134.